# FloPE: Flower Pose Estimation for Precision Pollination

### v1

@@ I INTRODUCTION
  • Codebase, data, models, and tools are freely available at https://github.com/wvu-irl/flope.
@@ II-B Object Pose Estimation
Direct approaches like PoseNet  [34] and  [28] regress camera pose from a single image by minimizing the Euclidean norm between the estimated and ground truth camera translations and rotations. A similar strategy is employed by [35] to regress object pose from a single image. However, the quaternions do not lie in the Euclidean manifold, but rather in the $SO(3)$ manifold. Hence, directly optimizing for Euclidean angles or quaternions will cause a sub-optimal learning path due to discontinuity and double cover [36, 37]. To address this issue, [38] learns rotational features in $\mathbb{R}^{6}$, and projects them to $SO(3)$ using Gram-Schmidt orthonormalization, providing continuous representation space. Another approach proposed by [39] suggests to learn rotational features in $\mathbb{R}^{9}$ and project it to $3\times 3$ rotational matrix in $SO(3)$ using singular value decomposition (SVD).
@@ III-A Data Generation
$g_{i}(\xi)=\sigma(\alpha_{i})\exp(-\frac{1}{2}(\xi-x_{i})^{T}\Sigma_{i}^{-1}(\xi-x_{i}))\vspace{-2mm}$ (1)
@@
We developed a 3D pose annotator tool that significantly simplifies annotating flower poses within the 3DGS point cloud. With this, a complete 3DGS plant model with around $20-25$ flowers can be annotated in a few minutes. The annotation process relies on human judgment to determine each flower’s orientation, which introduces some degree of error, but remains a practical and reliable approach for obtaining ground-truth data. Using the annotated 3D model, flower poses are projected onto the rendered image $I^{G}$ from multiple viewpoints to obtain the ground-truth orientation, $R^{\text{gt}}$, which is then utilized for training the pose estimation module.
@@ III-C Flower Pose Estimation
PoseNet [34] is utilized as the orientation estimation model $\mathcal{O}(I[b])$ to directly regress the rotational features $\rho\in\mathbb{R}^{9}$. The rotational features $\rho$ are reshaped to a $3\times 3$ matrix and projected to a rotational flower representation $R^{f}\in SO(3)$ using Singular Value Decomposition. Here, [.][.][ . ] denotes a cropping operation, where bounding boxes $b$ crop out patches from $I$, which are fed into $\mathcal{O}$. The choice of using $\mathbb{R}^{9}+\text{SVD}$ based rotational representation is vital for the learning process of the $\mathcal{O}$. This provides a continuous representation space without ambiguity (e.g., double cover), ensuring stable and nonfluctuating gradients during training. The model $\mathcal{O}$ is trained by minimizing $||R^{f}-R^{\text{gt}}||$, where ||.||||.||| | . | | represents the L2 norm. To account for the radial symmetry of the flowers, yaw angles are nullified before computing the loss. The camera intrinsic parameters $K$ and depth image $D$ are then used to map the 2D pixel coordinates $u$ to flower 3D coordinates $x^{f}$ as shown in Equation 2.
@@ III-D Pose Refinement
Since $x^{f}$ belongs to $\mathbb{R}^{3}$, filtering is simple with smooth interpolation between the predicted and measured state. However, orientation in $R^{f}\in SO(3)$ suffers from discontinuities and singularities when represented using Euler angles or quaternions, making direct filtering unstable; thus, it is converted to the $\mathbb{R}^{9}+\text{SVD}$ representation, with smooth interpolation [39]. Specifically, the rotational state and measurements are represented in $\mathbb{R}^{9}$ by flattening the $3\times 3$ rotation matrix $R^{f}$. The Kalman update step is then performed in this space. After each update, the resulting state is projected back onto $SO(3)$ using SVD before proceeding to the next Kalman update step. This ensures that the rotational state remains within $SO(3)$ while being smoothly updated with new rotational measurements.
@@ IV-A Flower Pose Estimation Accuracy
Evaluating predicted flower poses on a real plant is challenging due to the absence of reliable ground truth data. To address this, we generate a test dataset using a separate 3DGS plant model with accurately labeled flower pose annotations, with a different flower distribution than the training data. The evaluation metrics, summarized in Table I and visualized in Fig. 4, fall into three key categories: 2D metrics, which assess the accuracy of flower detection in the image plane; 3D metrics, which evaluate the precision of 6-degree-of-freedom (6-DoF) flower pose estimation; and speed metrics, which measure the computational performance of the models across different hardware platforms. We report these metrics for both the teacher and student models, as well as FloPE, our complete implementation.
Given an input image size of 1280x720, both models achieve a low detection error of approximately 8 pixels. We define a detection as successful if the error is within 20 pixels, which is used to compute the detection success rate. For pose estimation, we achieved approximately $3$ cm, $30\degree$ error before the refinement step. Since the detection error is minimal, most of the translational error arises from inaccuracies in depth estimation provided by the RealSense D405 camera. We found that depth measurements of this camera become unreliable outside its optimal operational range of 7–50 cm. The higher rotational error is likely due to ambiguities in the flower’s visual appearance from certain viewpoints. Accurately determining a flower’s precise orientation within a few degrees is inherently challenging, even for human observers when relying on a single viewpoint. To improve annotation accuracy, our annotation tool provides the ability to rotate and scale the model as needed, reducing pone annotation error.

### v2

@@ I INTRODUCTION
  • Codebase, data, models, and tools are freely available at wvu-irl.github.io/flope-irl.
@@ II-B Object Pose Estimation
Direct approaches like PoseNet  [34] and  [28] regress camera pose from a single image by minimizing the Euclidean norm between the estimated and ground truth camera translations and rotations. A similar strategy is employed by [35] to regress object pose from a single image. However, the quaternions do not lie in the Euclidean manifold, but rather in the $SO(3)$ manifold. Hence, directly optimizing for Euclidean angles or quaternions will cause a sub-optimal learning path due to discontinuity and double cover [36, 37]. To address this issue, [38] learns rotational features in $\mathbb{R}^{6}$, and projects them to $SO(3)$ using Gram-Schmidt orthonormalization, providing continuous representation space. Another approach proposed by [39] suggests to learn rotational features in $\mathbb{R}^{9}$ and project it to $\times 3$ 3 rotational matrix in $SO(3)$ using singular value decomposition (SVD).
@@ III-A Data Generation
$g_{i}(\xi)=\sigma(\alpha_{i})\exp(-\frac{1}{2}(\xi-x_{i})^{T}\Sigma_{i}^{-1}(\xi-x_{i}))\vskip-5.69054pt$ (1)
@@
We developed a 3D pose annotator tool that significantly simplifies annotating flower poses within the 3DGS point cloud. With this, a complete 3DGS plant model with around 202520-2520 - 25 flowers can be annotated in a few minutes. The annotation process relies on human judgment to determine each flower’s orientation, which introduces some degree of error, but remains a practical and reliable approach for obtaining ground-truth data. Using the annotated 3D model, flower poses are projected onto the rendered image $I^{G}$ from multiple viewpoints to obtain the ground-truth orientation, $R^{\text{gt}}$, which is then utilized for training the pose estimation module.
@@ III-C Flower Pose Estimation
PoseNet [34] is utilized as the orientation estimation model $\mathcal{O}(I[b])$ to directly regress the rotational features $\rho\in\mathbb{R}^{9}$. The rotational features $\rho$ are reshaped to a $\times 3$ 3 matrix and projected to a rotational flower representation $R^{f}\in SO(3)$ using Singular Value Decomposition. Here, [.][.][ . ] denotes a cropping operation, where bounding boxes $b$ crop out patches from $I$, which are fed into $\mathcal{O}$. The choice of using $\mathbb{R}^{9}+\text{SVD}$ based rotational representation is vital for the learning process of the $\mathcal{O}$. This provides a continuous representation space without ambiguity (e.g., double cover), ensuring stable and nonfluctuating gradients during training. The model $\mathcal{O}$ is trained by minimizing $||R^{f}-R^{\text{gt}}||$, where ||.||||.||| | . | | represents the L2 norm. To account for the radial symmetry of the flowers, yaw angles are nullified before computing the loss. The camera intrinsic parameters $K$ and depth image $D$ are then used to map the 2D pixel coordinates $u$ to flower 3D coordinates $x^{f}$ as shown in Equation 2.
@@ III-D Pose Refinement
Since $x^{f}$ belongs to $\mathbb{R}^{3}$, filtering is simple with smooth interpolation between the predicted and measured state. However, orientation in $R^{f}\in SO(3)$ suffers from discontinuities and singularities when represented using Euler angles or quaternions, making direct filtering unstable; thus, it is converted to the $\mathbb{R}^{9}+\text{SVD}$ representation, with smooth interpolation [39]. Specifically, the rotational state and measurements are represented in $\mathbb{R}^{9}$ by flattening the $\times 3$ 3 rotation matrix $R^{f}$. The Kalman update step is then performed in this space. After each update, the resulting state is projected back onto $SO(3)$ using SVD before proceeding to the next Kalman update step. This ensures that the rotational state remains within $SO(3)$ while being smoothly updated with new rotational measurements.
@@ IV-A Flower Pose Estimation Accuracy
Evaluating predicted flower poses on a real plant is challenging due to the absence of reliable ground truth data. To address this, we generate a test dataset using a separate 3DGS plant model with accurately labeled flower pose annotations, with a different flower distribution than the training data. The evaluation metrics, summarized in Table I and visualized in Fig. 4, fall into three key categories: 2D metrics, which assess the accuracy of flower detection in the image plane; 3D metrics, which evaluate the precision of 6-degree-of-freedom (6-DoF) flower pose estimation; and speed metrics, which measure the computational performance of the models across different hardware platforms. Here, rotational error represents angle between direction of flower (i.e. blue z-axis in Figure 5) in predicted and ground-truth flower poses. We report these metrics for both the teacher and student models, as well as FloPE, our complete implementation.
Given an input image size of 1280x720, both models achieve a low detection error of approximately 8 pixels. We define a detection as successful if the error is within 20 pixels, which is used to compute the detection success rate. For pose estimation, we achieved approximately 333 cm, 30°30$\degree$30 ° error before the refinement step. Since the detection error is minimal, most of the translational error arises from inaccuracies in depth estimation provided by the RealSense D405 camera. We found that depth measurements of this camera become unreliable outside its optimal operational range of 7–50 cm. The higher rotational error is likely due to ambiguities in the flower’s visual appearance from certain viewpoints. Accurately determining a flower’s precise orientation within a few degrees is inherently challenging, even for human observers when relying on a single viewpoint. To improve annotation accuracy, our annotation tool provides the ability to rotate and scale the model as needed, reducing pone annotation error.